# LLM Jailbreak Attack versus Defense Techniques - A Comprehensive Study

### v1

@@ Abstract
LLM Jailbreak Attack versus Defense Techniques - A Comprehensive Study
[…]
@@ 1 Introduction
Despite the various jailbreak attack and defense methodologies, to the best of our knowledge, there remains a significant gap in the literature regarding comprehensive evaluations of how well the attack methodologies can perform against defended LLMs and how well defense mechanisms against jailbreak attacks.
@@ 2.1 LLM Jailbreak
Jailbreak attacks on LLMs involve crafting prompts that exploit the models to generate malicious content. Despite the potential for harm, such as generating instructions for fabricating explosives, LLMs typically refrain from producing such responses due to the incorporation of safeguards during their training. These measures include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Robustness via Additional Fine-Tuning (RAFT) (Dong et al., 2023), and Preference Optimized Ranking (PRO) (Song et al., 2023), which collectively ensure the model’s adherence to ethical guidelines.
@@
Schwinn et al. (2023) An approach parallel to Zou et al. (2023), but at the sentence level
@@
Du et al. (2023)* Using LLM intrinsic propensity to safety or not-aligned that is dependent on the previous prompts
@@ 2.2 Jailbreak Attack Techniques
To provide a structured overview of the strategies utilized to compromise LLMs, we categorize current attack techniques into three categories, reflecting their fundamental traits. The first category, Generative Techniques, includes attacks that are dynamically produced, eschewing predetermined plans. The second category, Template Techniques, comprises attacks conducted via pre-defined templates or modifications in the generation settings. The last category, Training Gaps Techniques, focuses on exploiting weaknesses due to insufficient safeguards in safe training practices, such as RLHF (Ouyang et al., 2022). The techniques employed in our study are elaborated in Table 1, highlighting the models chosen for evaluation within our framework.
@@
Perplexity Hu et al. (2023) Calculates token-level perplexity using a probabilistic graphical model and evaluates the likelihood of each token being part of a malicious suffix.
@@ 2.3 Jailbreak Defense Techniques
We further conduct thorough examination on the existing defense mechanisms, classifying them into three primary categories based on their operational principles: Self-Processing Defenses, which rely exclusively on the LLM’s own capabilities; Additional Helper Defenses, which require the support of an auxiliary LLM for verification purposes; and Input Permutation Defenses, which manipulate the input prompt to detect and counteract malicious requests aimed at exploiting gradient-based vulnerabilities. An overview of these defense mechanisms is presented in Table 2.
@@ 3 Study Design
Baseline Selection Our methodology selection criteria were predicated on the method’s popularity and accessibility. For RQ1, our analysis covers nine attack methods, divided into five generative (AutoDAN (Liu et al., 2023a), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), GPTFuzz (Yu et al., 2023), GCG (Zou et al., 2023)) and four template-based approaches (Jailbroken (Wei et al., 2023), 77 Templates from existing study (Liu et al., 2023b), Deep Inception (Li et al., 2023a), Parameters (Huang et al., 2024)). The distinction between generative and template-based strategies aids in comprehensive coverage of attack methodologies.
For RQ2, we examine four defensive strategies: Bergeron (Pisano et al., 2023) for additional helper methods; RALLM (Cao et al., 2023) and SmoothLLM (Robey et al., 2023) for input permutation techniques; and Baseline (Jain et al., 2023) for perplexity analysis. Notable open-source projects, Aegis (Automorphic, 2023) and LLMguard (ProtectAI, 2023), alongside the OpenAI Moderation API (OpenAI, 2023), are also evaluated for their defense efficacy. Limitations such as Rain’s (Ouyang et al., 2022) lengthy processing and Certifying-llm’s (Kumar et al., 2023) scalability issues are acknowledged to refine our selection criteria.
@@ 3.1 LLMs under Test
In our research, we focus on evaluating three distinguished models: Llama-2-7b (Hugging Face, 2023a), Vicuna-v1.5-7b (Hugging Face, 2023b), and GPT-3.5-Turbo-1106 (OpenAI, 2023b). These models were chosen due to their prevalent use in security-related research, encompassing both attack simulations and the development of defensive strategies. The decision to omit GPT-4 from our evaluation stems from its significant operational requirements. Preliminary assessments of GPT-3.5-Turbo indicated extensive query processing times, reaching a total of 79,314, in addition to the economic implications linked to GPT-4’s token pricing—set at \$0.01 per 1,000 tokens (OpenAI, 2023a). These constraints make the inclusion of GPT-4 in a comparative analysis economically prohibitive.
@@ 3.2 Experimental Configuration
Our experimental framework utilized two NVIDIA RTX 6000 Ada GPUs, each outfitted with 48 GB of RAM. We aligned our testing parameters with those identified as optimal in the relevant literature, defaulting to the original repositories’ settings in the absence of specific recommendations. To address Research Question 1 (RQ1) and ensure consistency across different attack methodologies, each query was executed five times to minimize variability. For the evaluation involving generative models, we capped the process at a maximum of 75 iterations for each query, defining an iteration as a single algorithmic step.
@@ 3.3 Benchmark Construction
We leveraged the benchmark framework proposed by Liu et al. (Liu et al., 2023b). This benchmark is distinguished by its rigorous focus on policy compliance within the context of malicious content detection. In an effort to enhance the robustness of our evaluation, we expanded the original dataset to include 60 malicious queries, effectively doubling its size. This augmentation was achieved through meticulous manual curation and integrating selected examples from AdvBench (Zou et al., 2023). Our approach to dataset expansion adhered strictly to the categorization and selection criteria established in previous studies, ensuring both the consistency and the relevance of the enhanced dataset for comprehensive evaluation.
@@ 3.4 Result Labeling
In our study, we employed both automated and manual labeling strategies to categorize the responses gathered from our evaluation process. With a dataset encompassing 199,427 responses, the task of identifying malicious content posed significant challenges due to the impracticality of manual review and the absence of standardized evaluation methods for malicious responses—a gap in current research.
Existing strategies for addressing this issue vary. Zou et al. (Zou et al., 2023) utilized a set of common refusal patterns, such as "I am sorry" and "I cannot", to automate the identification of non-compliant responses. Yu et al. (Yu et al., 2023) and Huang et al. (Huang et al., 2024) focused on enhancing machine learning models, specifically RoBERTa and BERT-BASE-CASED models, respectively. Additionally, approaches leveraging GPT-4 for attack analysis were explored by Chao et al. (Chao et al., 2023) and Mehrotra et al. (Mehrotra et al., 2023).
To ensure a rigorous analysis of our dataset, we selected a random sample of 1,068 responses, aiming for a 95% confidence level with a 3% margin of error. This sample underwent manual annotation to assess each response’s relevance to the inquiry and its alignment with providing a malicious answer. An example of this annotation process is documented in Table 6, and a comparative analysis of the effectiveness of different models is provided in Table 5.
Reflecting on the reported efficacy of the fine-tuned RoBERTa model by Yu et al. (Yu et al., 2023), we chose to refine this model further utilizing our manually annotated dataset, accessible on HuggingFace (fine tuned, 2024). The fine-tuning protocol involved a batch size of 5, three training epochs, a learning rate of $2\times 10^{-5}$, application of the Adam optimizer, and a linear rate decay complemented by a warm-up phase covering 10% of the training duration. Post-labeling, an additional round of random sampling was conducted for manual verification to ascertain the accuracy and reliability of our findings.
Evaluation Metric For RQ1, we use two metrics.
This dual metric approach ensures a comprehensive evaluation of both the attack’s impact and its operational feasibility. First, Attack Success Rate (ASR): defined as the ratio of successfully compromised questions c to the total number of questions n, ASR measures the effectiveness of an attack.
@@
The second metric, Benign Passing Rate (BPR), assesses the proportion of non-malicious inputs s that successfully navigate through the defense filter relative to the total number of inputs t.
$BPR=\frac{s}{t}.$ (4)
@@ 4 RQ1: Effectiveness of Jailbreak Attack
The effectiveness of attack strategies on the selected LLMs under test is systematically presented in Tables 67, and 8. These tables enumerate the number of queries that were effectively bypassed under various categories, corresponding to the detailed breakdowns provided in Tables 910, and 11. To offer a clearer comparative analysis of model performance, we consolidated these metrics into a scatter plot depicted in Figure 3. In this visualization, models demonstrating optimal performance are positioned nearer to the scatter plot’s upper right quadrant, signifying superior ASR and Efficiency.
Evaluation results reveal that using 78 templates, Jailbroken, and GPTFuzz strategies yield superior results in circumventing the security measures of GPT-3.5-turbo and Vicuna. Conversely, for LLaMA, strategies such as Jailbroken, Parameter, and 78 templates demonstrated the highest effectiveness. This prevalence of template-based approaches highlights their efficiency, primarily due to the intricate design of their prompts. The most successful templates from these strategies are detailed in Table 16, which will be added subsequently.
In the realm of generative strategies, GPTFuzz, Pair, and Tap emerged as the top performers. Moreover, it was noted that LLaMA presents a noteworthy challenge for jailbreaking compared to Vicuna. Specifically, Vicuna showed a quicker reduction in loss when employing the AutoDan method and achieved jailbreak at a higher loss threshold. In contrast, LLaMA exhibited a slower loss reduction rate and, despite achieving a lower loss than Vicuna, did not succumb to jailbreaking, as illustrated in Figure 2. Additionally, our study into the categories of questions that were successfully jailbroken indicates that queries related to harmful content and illegal activities are the most challenging to address across all tested models.
Figure 2: Loss of a random question
@@ 5 RQ2: Effectiveness of Jailbreak defense
Our study meticulously evaluates defense mechanisms against malicious queries as well as the handling of benign questions. The outcomes of this evaluation are systematically tabulated in Tables 1413, and 12. These results are further visualized in Figure 4, where the optimal defense strategies are identified by their proximity to the upper left corner of the plot, signifying lower DPR and higher BPR. Our findings reveal that, apart from the Bergeron method, the efficacy of the current defense strategies remains largely inadequate. Additionally, our comparative analysis of the quality of benign responses generated through three innovative methodologies disclosed minimal variance among them, as elaborated in Table 15.
@@ 6.1 Comparative Performance of White-Box and Black-Box Attacks
To further understand the influence of loss metrics on a model’s vulnerability to jailbreaking, we conducted a targeted experiment. A question was randomly selected from our dataset, and the experiment’s findings are visually represented in Figure 2. The experiment showed that Vicuna began the process with a higher initial loss but saw a significant reduction in loss, stabilizing after 12 steps and five successful jailbreak attempts. However, it maintained a higher final loss compared to LLaMa. In contrast, LLaMa started with a lower initial loss and demonstrated a slower reduction in loss over time, ultimately failing to jailbreak the question despite exhibiting a significantly lower final loss than Vicuna. These outcomes suggest that LLaMa’s foundational safety training plays a pivotal role in its enhanced defense against jailbreak attempts. It implies that integrating advanced safety training protocols into developing open-source models could markedly reduce the efficacy of white-box attacks, thereby enhancing their security posture.
@@ 6.3 Enhancing Defense Mechanisms Against Diverse Malicious Queries
Our research underscores the imperative need for robust defense mechanisms to be rigorously evaluated against a broad spectrum of malicious queries. The development of a uniform and standardized evaluation methodology emerges as a pivotal challenge within the realm of defense strategies. Present techniques, as outlined by Kumar et al. (Kumar et al., 2023), Robey et al. (Robey et al., 2023), and Cao et al. (Cao et al., 2023), rely on predefined sets that frequently misclassify benign responses, such as "I’m sorry, I do not understand," as malicious. This issue of misclassification becomes even more pronounced when incorporating models like GPT-4 with these sets, as GPT-4’s limitations in generating contextually accurate responses can lead to benign inputs being wrongly identified as malicious, as illustrated in Figure 5.
Moreover, our analysis reveals that the efficacy of existing defense strategies is generally insufficient. While the Bergeron method shows promise, its effectiveness is limited by its dependence on LLMs and the use of predefined prompts for processing and verification, which not only undermines its reliability but also entails substantial operational costs. Proposals by Kumar et al., Robey et al., and Cao et al. for identifying malicious prompts through suffix addition demonstrate potential but struggle with scalability issues, particularly when dealing with complex natural language inputs.
Given these observations, there is a critical and pressing need for further research into more advanced evaluation frameworks and formulating more effective defense strategies. Such efforts should aim to circumvent the current limitations by ensuring reliable differentiation between malicious and benign inputs across varying contexts and increasing the scalability of defense mechanisms to accommodate the complexities inherent in natural language processing.

### v2

# A Comprehensive Study of Jailbreak Attack versus Defense for Large Language Models

@@ Abstract
Warning:This paper contains unsafe model responses.
@@
A Comprehensive Study of Jailbreak Attack versus Defense for Large Language Models
Zihao Xu1,2,∗  Yi Liu3,†  Gelei Deng3,‡  Yuekang Li1,§  Stjepan Picek2,¶ 1University of New South Wales, Australia 2Delft University of Technology, The Netherlands 3Nanyang Technological University, Singapore zhltroin@gmail.com, yi009@e.ntu.edu.sg, gelei.deng@ntu.edu.sg §yuekang.li@unsw.edu.au, S.Picek@tudelft.nl
@@ 1 Introduction
Despite the various jailbreak attack and defense methodologies, to the best of our knowledge, there remains a significant gap in the literature regarding comprehensive evaluations of how well the attack methodologies can perform against defended LLMs and how well defense mechanisms against jailbreak attacks. While  Mazeika et al. (2024) and Zhou et al. (2024) explore various attack techniques, they did not evaluate those on defense techniques, and vice versa.
@@ 2 Background and Related Work
This study underscores the effectiveness of specific attack methodologies against various defense strategies and vice versa, filling a gap not addressed in contemporary literature (Mazeika et al., 2024; Zhou et al., 2024). These works primarily focus on evaluating various attack techniques against unprotected models, with the exception of initial safety training. Our research conducts the first comprehensive survey that evaluates the reciprocal impacts of both attack and defense techniques.
@@ 2.1 LLM Jailbreak
Jailbreak attacks on LLMs involve crafting prompts that exploit the models to generate malicious content. Despite the potential for harm, such as generating instructions for fabricating explosives, LLMs typically refrain from producing such responses due to the incorporation of safeguards during their training. These measures include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Robustness via Additional Fine-Tuning (RAFT) (Dong et al., 2023), and Preference Optimized Ranking (PRO) (Song et al., 2023), which ensure the model’s adherence to ethical guidelines.
@@
In the subsequent subsection, we present a categorization of current attack and defense techniques. Additionally, we analyze the pros and cons of each category in various dimensions. Details can be found in A.1 and A.2. This analysis facilities a comprehensive understanding and substantiating our categorization approach.
@@
Schwinn et al. (2023) An approach parallel to Zou et al. (2023), but at the sentence level, and focus on optimizing the whole given suffix in continuous values.
@@
Du et al. (2023) Using LLM intrinsic propensity to safety or not-aligned that is dependent on the previous prompts
@@ 2.2 Jailbreak Attack Techniques
To provide a structured overview of the strategies utilized to compromise LLMs, we categorize current attack techniques into three categories, reflecting their fundamental traits. The first category, Generative Techniques, includes attacks that are dynamically produced, eschewing predetermined plans. The second category, Template Techniques, comprises attacks conducted via pre-defined templates or modifications in the generation settings. The last category, Training Gaps Techniques, focuses on exploiting weaknesses due to insufficient safeguards in safe training practices, such as RLHF (Ouyang et al., 2022). The techniques employed in our study are elaborated in Table 1, highlighting the methods chosen for evaluation within our framework.
@@
Hu et al. (2023) Calculates token-level perplexity using a probabilistic graphical model and evaluates the likelihood of each token being part of a malicious suffix.
@@ 2.3 Jailbreak Defense Techniques
We further conduct thorough examination on the existing defense mechanisms, classifying them into three categories based on their operational principles: Self-Processing Defenses, which rely exclusively on the LLM’s own capabilities; Additional Helper Defenses, which require the support of additional algorithms or auxiliary LLMs for verification purposes; and Input Permutation Defenses, which manipulate the input prompt and veritfy with the target LLMs mutliple times to detect and counteract malicious requests aimed at exploiting gradient-based vulnerabilities. An overview of these defense mechanisms is presented in Table 2.
@@ 3 Study Design
3.1 Baseline Selection
Our methodology selection criteria were predicated on the method’s popularity and accessibility to source code. For RQ1, our analysis covers nine attack techniques, divided into five generative (AutoDAN (Liu et al., 2023a), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), GPTFuzz (Yu et al., 2023), GCG (Optimize per prompt on a single model) (Zou et al., 2023)) and four template-based approaches (Jailbroken (Wei et al., 2023), 77 Templates from existing study (Liu et al., 2023b), Deep Inception (Li et al., 2023a), Parameters (Huang et al., 2024)). To elucidate the characteristics of the prompts used in attack techniques, we present an illustrative example in Figure 7.
For RQ2, we examine four defense techniques: Bergeron (Pisano et al., 2023) and Baseline (Jain et al., 2023) for additional helper methods; RALLM (Cao et al., 2023) and SmoothLLM (Robey et al., 2023) for input permutation techniques; Notable open-source projects, Aegis (Automorphic, 2023) and LLMguard (ProtectAI, 2023), alongside the OpenAI Moderation API (OpenAI, 2023), are also evaluated for their defense efficacy. Limitations such as Rain’s (Ouyang et al., 2022) extensively prolonged time-consuming processing and Certifying-llm’s (Kumar et al., 2023) scalability issues are considered to be excluded from our selection.
@@ 3.2 LLMs under Test
In our research, we focus on evaluating three distinguished models: Llama-2-7b (Hugging Face, 2023a), Vicuna-v1.5-7b (Hugging Face, 2023b), and GPT-3.5-Turbo-1106 (OpenAI, 2023b). These models were chosen due to their prevalent use in security-related research, encompassing both attack simulations and the development of defensive strategies. The decision to omit GPT-4 from our evaluation stems from its significant operational requirements. Preliminary evaluations of GPT-3.5-Turbo revealed an exceptionally high query count, totaling 79,314. When taking into account the economic ramifications associated with the token pricing of GPT-4, which is established at \$0.01 per 1,000 tokens (OpenAI, 2023a), this financial consideration renders the incorporation of GPT-4 into a comparative study economically challenging.
@@ 3.3 Experimental Configuration
Our experimental framework utilized two NVIDIA RTX 6000 Ada GPUs, each outfitted with 48 GB of RAM. We aligned our testing parameters with those identified as optimal in the relevant literature, defaulting to the original repositories’ settings in the absence of specific recommendations. To address RQ1 and ensure consistency across different attack methodologies, each query was executed 5 times to minimize variability. For the evaluation involving generative models, we capped the process at a maximum of 75 iterations for each query, defining an iteration as a single algorithmic step. However, in our empirical study of GCG with 18 questions that randomly and uniformly sampled from six categories, it suggests that GCG only on Llama requires a higher number of iterations to jailbreak most queries otherwise failure. In order to not be biased to GCG, we use the default 500 iterations on Llama model only. We provide a further discussion in Section 6.1
@@ 3.4 Benchmark Construction
We leveraged the benchmark framework proposed by Liu et al. (Liu et al., 2023b). This benchmark is distinguished by its rigorous focus on policy compliance to OpenAI categories(OpenAI, 2023) within the context of malicious content detection. In an effort to enhance the robustness of our evaluation, we expanded the original dataset to include 60 malicious queries, effectively doubling its size. This augmentation was achieved through meticulous manual curation and integrating selected examples from AdvBench (Zou et al., 2023). Our approach to dataset expansion adhered strictly to the categorization and selection criteria established in previous studies, ensuring both the consistency and the relevance of the enhanced dataset for comprehensive evaluation.
@@ 3.5 Result Labeling
In our study, we employed both automated and manual labeling strategies to categorize the responses gathered from our evaluation process, details can be found in Appendix A.3.
3.6 Evaluation Metric
For RQ1, we use two metrics. This dual metric approach ensures a comprehensive evaluation of both the attack’s impact and its operational feasibility. First, Attack Success Rate (ASR): defined as the ratio of successfully compromised questions c to the total number of questions n, ASR measures the effectiveness of an attack.
@@ 3.6 Evaluation Metric
The second metric, Benign Success Rate (BSR), assesses the proportion of non-malicious inputs s that successfully navigate through the defense filter relative to the total number of inputs t.
$BSR=\frac{s}{t}.$ (4)
@@ 4 RQ1: Effectiveness of Jailbreak Attack
The effectiveness of attack strategies on the selected LLMs under test is systematically presented in Tables 67, and 8. To offer a clearer comparative analysis of model performance, we consolidated these metrics into a scatter plot depicted in Figure 2. In this visualization, models demonstrating optimal performance are positioned nearer to the scatter plot’s upper right quadrant, signifying superior ASR and Efficiency.
Evaluation results reveal that using 78 templates, Jailbroken, and GPTFuzz strategies yield superior results in circumventing the security measures of GPT-3.5-turbo and Vicuna. Conversely, for LLaMA, strategies such as Jailbroken, Parameter, and 78 templates demonstrated the highest effectiveness. This prevalence of template-based approaches highlights their efficiency, primarily due to the intricate design of their prompts. The most successful five templates from these strategies are listed in Table 16.
In the realm of generative strategies, GPTFuzz, Pair, and Tap emerged as the top performers. Moreover, it was noted that LLaMA presents a noteworthy challenge for jailbreaking compared to Vicuna. We will discuss this in Section 6.1. Additionally, our study into the categories of questions that were successfully jailbroken indicates that queries related to unlawful practice, harmful content and illegal activities are the most challenging to address across all tested models. Details can be found in Table 9, 11 and 10.
@@ 5 RQ2: Effectiveness of Jailbreak defense
Our study meticulously evaluates defense mechanisms against malicious queries as well as the handling of benign questions. The outcomes of this evaluation are systematically tabulated in Tables 1213, and 14. These results are further visualized in Figure 3, where the optimal defense strategies are identified by their proximity to the upper left corner of the plot, signifying lower DPR and higher BSR. Our findings reveal that, apart from the Bergeron method, the efficacy of the current defense strategies remains largely inadequate. Additionally, our comparative analysis of the quality of benign responses generated through three innovative methodologies disclosed minor variance among them, as elaborated in Table 15.
@@ 6.1 Comparative Performance of White-Box and Black-Box Attacks
To further understand the influence of loss metrics on a model’s vulnerability to jailbreaking, we conducted a targeted experiment. A question was randomly selected from our dataset, and the experiment’s findings are visually represented in Figure 4. The experiment showed that Vicuna began the process with a higher initial loss but saw a significant reduction in loss, stabilizing after 12 steps and five successful jailbreak attempts. However, it maintained a higher final loss compared to LLaMa. In contrast, LLaMa started with a lower initial loss and demonstrated a slower reduction in loss over time, ultimately failing to jailbreak the question within 75 iteration steps despite exhibiting a significantly lower final loss than Vicuna. These outcomes suggest that LLaMa’s foundational safety training plays a pivotal role in its enhanced defense against jailbreak attempts. It implies that integrating advanced safety training protocols into developing open-source models could markedly reduce the efficacy of white-box attacks, thereby enhancing their security posture.
Figure 4: Loss of a random question
@@ 6.3 Enhancing Defense Mechanisms Against Diverse Malicious Queries
Our research underscores the imperative need for robust defense mechanisms to be rigorously evaluated against a broad spectrum of malicious queries. The development of a uniform and standardized evaluation methodology emerges as a primary challenge within the realm of defense strategies. Present techniques, as outlined by Kumar et al. (2023),  Robey et al. (2023), and Cao et al. (2023), rely on predefined sets that frequently misclassify benign responses, such as "I’m sorry, I do not understand," as malicious. This issue of misclassification becomes even more pronounced when incorporating models like GPT-4 with these sets, as GPT-4’s limitations in generating contextually accurate responses can lead to benign inputs being wrongly identified as malicious, as illustrated in Figure 8.
The lack of a uniform evaluation methodology is also reflected by the challenge, Detection Constraints. Although (OpenAI, 2023) categorizes several forms of toxic content, the diversity in expressions and responses can overwhelm sophisticated classifiers, as evidenced by the performance of the ModerationAPI. Moreover, the efficacy of detection models is inherently limited by their model characteristics, such as the number of parameters, model structure and training dataset, etc.
Moreover, our analysis reveals that another challenge, Cost Concerns. While the Bergeron method shows promise, its effectiveness is limited by its dependence on LLMs and the use of predefined prompts for processing and verification, which not only undermines its reliability but also entails substantial operational costs.
Another concern is Latency Issues. In our experiments, we observed that references RALLM (Cao et al., 2023) and Smooth-LLM (Robey et al., 2023) exhibit significantly longer processing times for queries and struggle with scalability issues, particularly when dealing with complex natural language inputs. This issue is attributable to their design approach, which involves repeatedly checking a single prompt variation, consequently prolonging the defense time. A similar issue was also noted in our preliminary experiments with RAIN (Li et al., 2023c).
Given these observations, there is a critical and pressing need for further research into more advanced evaluation frameworks and formulating more effective defense strategies. Such efforts should aim to circumvent the current challenges by ensuring reliable differentiation between malicious and benign inputs across varying contexts and increasing the scalability of defense mechanisms to accommodate the complexities inherent in natural language processing.